OpenStudy (anonymous):

f(x) = 2x^(4) − 4x^(2) - 48. Find both of the real zeros of the polynomial function.

OpenStudy (anonymous):

solve \[2z^2-4z-48=0\] then replace \[z=x^2\] and solve that one

OpenStudy (anonymous):

I already did that but my answers don't make sense.

OpenStudy (anonymous):

you get \[2(z-6)(z+4)=0\] \[z=6,z=-4\] so \[x^2=6,x^2=-4\] the second one is not possible, so only two answers that make sense are \[x^2=6\] \[x=\pm\sqrt{6}\]